# Convolutional Neural Network NLP Natural Language Processing

### v1

@@ 5 Conclusion and Future Work
Table A.1: Performance for different token mixers $T$ across datasets. Each stage of the MetaFormer utilizes the given token mixer and was trained from scratch. Due to training instabilities within the implementation of local attention, we had to adapt the training routine (cf. Sec. 4.1.4). Please refer to markers ($\dagger$) in Tab. 2 metarow $[T,T,T,T]$ to identify those results.
@@
Table A.2: Performance for different token mixers $T$ across datasets. The MetaFormer utilizes ImageNet-pretrained weights for the channel-MLPs and only exchanges token mixers in the last two stages, keeping the original pooling at the first two stages ($[P,P,T,T]$).
@@
Table A.3: Performance (AUC and F1) for different token mixers $T$ across datasets. The MetaFormer utilizes pooling at the first two stages and the given token mixer at the last two stages ($[P,P,T,T]$). These results provide the base for the calculation of the percental improvement shown in Fig. 3(b).
@@
Table A.4: Detailed results by our ranking schema (cf. Sec. 3.3) for each token mixer across datasets. Each entry shows a tuple: the first element is the absolute number of pairwise wins against other token mixers, and the second is the normalized rank in $[0.1,1]$. Global ranks are obtained as the geometric mean of normalized ranks across datasets.
@@
Table A.5: Detailed results by our ranking schema (cf. Sec. 3.3) for each token mixer across datasets. Each entry shows a tuple: the first element is the absolute number of pairwise wins against other token mixers, and the second is the normalized rank in $[0.1,1]$. Global ranks are obtained as the geometric mean of normalized ranks across datasets.
@@
pooling 3 (4.0, 0.55) (2.0, 0.45) (6.0, 0.92) 0.608
5 (3.0, 0.45) (2.0, 0.45) (2.0, 0.55) 0.478
7 (1.0, 0.27) (2.0, 0.45) (2.0, 0.55) 0.406
9 0.6019
conv 3 (8.0, 0.69) (5.0, 0.72) (10.0, 1.0) 0.793
5 (10.0, 1.0) (6.0, 0.79) (1.0, 0.41) 0.687
7 (9.0, 0.86) (10.0, 1.0) (1.0, 0.41) 0.707
9 (3.0, 0.69) 0.5689
grouped conv 3 (9.0, 0.86) (2.0, 0.45) (2.0, 0.55) 0.596
5 (8.0, 0.69) (9.0, 0.93) (5.0, 0.83) 0.811
7 (9.0, 0.86) (8.0, 0.86) (4.0, 0.77) 0.832
9 (3.0, 0.69)
local attn 3 (0.0, 0.1) (2.0, 0.45) (0.0, 0.21) 0.212
5 (4.0, 0.55) (2.0, 0.45) (0.0, 0.21) 0.374
7 (1.0, 0.27) (2.0, 0.45) (0.0, 0.21) 0.296
9 (0.0, 0.21)
global attn - (1.0, 0.27) (0.0, 0.13) DNS 0.192
identity 1 (1.0, 0.27) (0.0, 0.13) (0.0, 0.21) 0.198

### v2

@@ Appendix A Overview of Appendix
Table 4: Performance for different token mixers $T$ across datasets. Each stage of the MetaFormer utilizes the given token mixer and was trained from scratch. Due to training instabilities within the implementation of local attention, we had to adapt the training routine (cf. Sec. 5.1.4). Please refer to markers ($\dagger$) in Tab. 2 metarow $[T,T,T,T]$ to identify those results.
@@
Table 5: Performance for different token mixers $T$ across datasets. The MetaFormer utilizes ImageNet-pretrained weights for the channel-MLPs and only exchanges token mixers in the last two stages, keeping the original pooling at the first two stages ($[P,P,T,T]$).
@@
Table 6: Performance (AUC and F1) for different token mixers $T$ across datasets. The MetaFormer utilizes pooling at the first two stages and the given token mixer at the last two stages ($[P,P,T,T]$). These results provide the base for the calculation of the percental improvement shown in Fig. 4(b).
@@
Table 7: Detailed results by our ranking schema (cf. Sec. 4.3) for each token mixer across datasets. Each entry shows a tuple: the first element is the absolute number of pairwise wins against other token mixers, and the second is the normalized rank in $[0.1,1]$. Global ranks are obtained as the geometric mean of normalized ranks across datasets.
@@
Table 8: Detailed results by our ranking schema (cf. Sec. 4.3) for each token mixer across datasets. Each entry shows a tuple: the first element is the absolute number of pairwise wins against other token mixers, and the second is the normalized rank in $[0.1,1]$. Global ranks are obtained as the geometric mean of normalized ranks across datasets.
@@
pooling 3 (4, 0.55) (2, 0.45) (6, 0.92) 0.608
5 (3, 0.45) (2, 0.45) (2, 0.55) 0.478
7 (1, 0.27) (2, 0.45) (2, 0.55) 0.406
9 (6, 0.92)
conv 3 (8, 0.69) (5, 0.72) (10, 1.0) 0.793
5 (10, 1.0) (6, 0.79) (1, 0.41) 0.687
7 (9, 0.86) (10, 1.0) (1, 0.41) 0.707
9 (3, 0.69)
grouped conv 3 (9, 0.86) (2, 0.45) (2, 0.55) 0.596
5 (8, 0.69) (9, 0.93) (5, 0.83) 0.811
7 (9, 0.86) (8, 0.86) (4, 0.77) 0.832
9 (3, 0.69)
local attn 3 (0, 0.1) (2, 0.45) (0, 0.21) 0.212
5 (4, 0.55) (2, 0.45) (0, 0.21) 0.374
7 (1, 0.27) (2, 0.45) (0, 0.21) 0.296
9 (0, 0.21)
global attn - (1, 0.27) (0, 0.13) DNS 0.192
identity 1 (1, 0.27) (0, 0.13) (0, 0.21) 0.198